# PaRCE: Probabilistic and Reconstruction-Based Competency Estimation for Safe Navigation Under Perception Uncertainty

### v1

@@ Abstract
Perception-based navigation systems are useful for unmanned ground vehicle (UGV) navigation in complex terrains, where traditional depth-based navigation schemes are insufficient. However, these data-driven methods are highly dependent on their training data and can fail in surprising and dramatic ways with little warning. To ensure the safety of the vehicle and the surrounding environment, it is imperative that the navigation system is able to recognize the predictive uncertainty of the perception model and respond safely and effectively in the face of uncertainty. In an effort to enable safe navigation under perception uncertainty, we develop a probabilistic and reconstruction-based competency estimation (PaRCE) method to estimate the model’s level of familiarity with an input image as a whole and with specific regions in the image. We find that the overall competency score can correctly predict correctly classified, misclassified, and out-of-distribution (OOD) samples. We also confirm that the regional competency maps can accurately distinguish between familiar and unfamiliar regions across images. We then use this competency information to develop a planning and control scheme that enables effective navigation while maintaining a low probability of error. We find that the competency-aware scheme greatly reduces the number of collisions with unfamiliar obstacles, compared to a baseline controller with no competency awareness. Furthermore, the regional competency information is very valuable in enabling efficient navigation.
@@ I INTRODUCTION
Figure 1: We develop a probabilistic reconstruction-based method for estimating the overall and regional competency of a perception model for a given input image. This competency score and competency map is used to design a sampling-based path planning algorithm that evaluates the minimum competency associated with sampled paths, removes trajectories that fall below the minimum probability threshold, and selects the best path towards the goal. An LQR-based controller is then used to track the planned path.
@@
To address these limitations, we develop a competency-aware perception, planning, and control framework to enable safe navigation in complex environments under DNN-based perception model uncertainty (Figure 1). In particular, (1) we develop an overall competency score that represents the probability that the model’s prediction is correct for a given image. (2) We propose a regional competency method that estimates the probability that regions in the environment are unfamiliar to the model. (3) We generate a method for evaluating the probability of error associated with a proposed control sequence based on the regional competency map. (4) We develop a navigation scheme that maintains efficient performance, while ensuring the probability of error remains below a user-defined error threshold. (5) We evaluate this control scheme in a challenging lunar environment with unfamiliar obstacles and demonstrate the ability of our navigation framework to successfully and efficiently navigate to goal regions, while maintaining a low probability of error.
@@ II BACKGROUND & RELATED WORK
Field robots used in search and rescue, planetary exploration, and agricultural tasks, must navigate in unstructured environments, which lack clearly viable paths or helpful landmarks [1]. In such environments, perception provides the necessary information to make the vehicle aware of the surrounding environment. Within this context, most work has focused on DNN-based methods to analyze the capability of a UGV to stably reach a terrain region and determine appropriate control actions [1]. Without properly considering the uncertainty surrounding the predictions of these DNN-based perception models, vehicles that rely on these models can face unsafe situations and system failures. This has led to a branch of uncertainty-aware navigation focused on dealing with perception model prediction uncertainty.
@@ II-A Uncertainty Quantification (UQ)
Predictive uncertainty generally encapsulates both data/aleatoric uncertainty, which arises from complexities of the data (i.e., noise, class overlap, etc.), and model/epistemic uncertainty, which reflects the ability of the perception model to capture the true underlying model of the data [8]. The modeling of these uncertainties can generally be divided into methods based on (1) Bayesian neural networks (BNNs) that extract uncertainty as a statistical measure over the output of a BNN [9, 10], (2) deterministic neural networks (NNs), often relying on Monte Carlo (MC) dropout as approximate Bayesian inference [11], and (3) ensembles of NNs that combine the predictions of multiple deterministic networks to form a probability density function [12]. While these methods address the typical overconfidence in NN predictions with better-calibrated confidence scores, they generally focus on predictions for in-distribution inputs, which come from the same distribution as the training data. These approaches are not sufficient in general to appropriately assign confidence scores for OOD inputs that differ significantly from those seen during training [4, 13].
@@ II-D Uncertainty-Aware Navigation
There are various approaches that seek to enable short-term, local control under perception model predictive uncertainty. Existing approaches tend to rely on one of the UQ methods discussed in Section II-A, using either BNNs [48, 49], MC dropout [50], or ensembles [51, 52, 53] to estimate uncertainty and inform the control scheme. As mentioned previously, while these methods can effectively estimate some aspects of predictive uncertainty, they are insufficient for quantifying uncertainty associated with OOD samples, which a UGV is bound to encounter when navigating previously unseen or dynamic environments. There has also been work that leverages uncertainty estimation approaches that relate more closely to the distance-based OOD detection methods [54] (using the estimation method developed in [55]) and classification-based anomaly detection methods [56] (using the method developed in [57]). While these approaches do not fully capture the predictive uncertainty of the perception models, they are better equipped to handle inputs that differ significantly from those seen during training.
We develop a reconstruction-based method for predictive uncertainty estimation with the end goal of probabilistically safe control of a UGV. Our approach for uncertainty estimation has several benefits compared to existing methods. (1) It captures multiple aspects of uncertainty and can effectively identify OOD or anomalous samples, unlike traditional UQ techniques. (2) In contrast to methods that simply seek to identify samples that differ significantly from the training set, our method is directly tied to the predictive uncertainty of the perception model used in a control task. Furthermore, rather than making a binary decision, as is typical in OOD and anomaly detection tasks, our method is probabilistic and can be incorporated into a probabilistically safe control scheme. (3) Finally, our method can characterize uncertainty at a regional level, rather than considering uncertainty only for the image as a whole. This offers a greater amount of model explainability and provides more flexibility in designing an uncertainty-aware planning and control scheme. We show how our method for predictive uncertainty estimation can be used to evaluate the probability of error associated with vehicle trajectories and to select a preferred trajectory based on the desired level of conservativeness. [^1] The code for reproducing our methods and results is available on GitHub: https://github.com/sarapohland/parce.git.
@@ IV-C Estimating Regional Model Competency
Suppose that, in addition to estimating the overall competency score for the image as a whole, we wish to estimate a regional competency score for each segmented region in the image. Now instead of letting $\bm{X}$ be the entire input image, $\bm{X}$ is a segmented region of the image. In this work, segmented regions are determined by the Felzenszwalb segmentation algorithm [58]. We follow roughly the same procedures to estimate the probability that each region in the input image came from the same distribution as the training samples. In this approach, rather than designing an autoecoder to reconstruct the input image, we design an image inpainting model to reconstruct a missing segment of the input image and measure the average reconstruction loss over the pixels corresponding to that image segment.
@@ V COMPETENCY-AWARE NAVIGATION
We use the overall model competency estimates (§IV-A) and regional competency estimates (§IV-C) to explore navigation schemes with varying levels of competency awareness. We develop a model of the vehicle dynamics (§V-A), use this model to design a competency-aware path planner (§V-B), and develop a path-tracking controller (§V-C).
@@ V-A Model of Vehicle Dynamics
Figure 4: Three examples of competency-aware planning. Column 1: Input images to perception model. Column 2: Probabilistic regional competency estimates. Blue indicates high competency (around 1.0) and red indicates low competency (around 0.0). Column 3: Minimum competency score associated with sampled trajectories. Column 4: Trajectories deemed safe by the competency-aware planner.
@@ V-B Competency-Aware Path Planning
We develop five planners that utilize different competency information and strategies for responding to low competency. (1) The planner that uses only overall competency selects the best trajectory in the same way as the baseline planner when overall competency is high. If the competency score falls below some probability threshold, the planner resorts to a safe action sequence, backing up slightly then turning. (2) The turning-based planner that utilizes only regional competency information checks if regions in the vicinity of the robot are associated with low perception model competency. If those regions fall below some probability threshold, the vehicle backs up slightly and turns away from low competency regions. (3) The turning-based planner that uses both overall and regional competency checks both measures of competency before resorting to a safe response. (4) The trajectory-based planner that uses only regional competency information maps sampled paths onto the regional competency map, determines the minimum competency value associated with each path, and removes those that fall below the threshold. The planner will then select the best path among those remaining. If there is no safe path, the vehicle will back up slightly and turn away from the low competency regions. (5) The trajectory-based planner that utilizes both overall and regional competency behaves similarly but only considers regional competency when overall competency is low. Three examples of this trajectory-based planner with full competency-awareness are shown in Figure 4.
@@ V-D Evaluation of Navigation Performance
We evaluate the navigation capability of the full perception, planning, and control pipeline (Figure 1) with varying levels of competency awareness and different responses to low competency. We compare (1) a baseline controller that uses no competency information, (2) a turning-based controller that uses only overall competency scores, (3) a turning-based controller that uses only regional competency maps, (4) a trajectory-based controller that uses only regional competency information, (5) a turning-based controller that uses both overall competency scores and regional competency maps, and (6) a trajectory-based controller that uses both overall and regional competency information.
@@
TABLE I: A comparison of (1) baseline controller with no competency awareness, (2) turning-based controller with only overall competency information, (3) turning-based controller with only regional competency maps, (4) trajectory-based controller with only regional competency information, (5) turning-based controller with overall competency and regional competency awareness, and (6) trajectory-based controller with both types of competency awareness.
@@ VI CONCLUSIONS
We aim to improve the safety of a UGV navigating according to a perception-based system in a complex environment with obstacles that are unfamiliar to the perception model. We believe that integrating information about the predictive uncertainty of the perception model into the navigation scheme will improve the safety and performance of the system. With this idea, we propose a novel overall competency score that can distinguish between correctly classified ID samples, misclassified ID samples, and OOD samples. We expand on this approach, to develop a regional competency map that can visualize regions in an image that are familiar to the perception model and those that are not. We then explore how this competency information can be integrated into planning and control scheme for UGV navigation.

### v3

# Competency-Aware Planning for Probabilistically Safe Navigation Under Perception Uncertainty

@@ Abstract
Perception-based navigation systems are useful for unmanned ground vehicle (UGV) navigation in complex terrains, where traditional depth-based navigation schemes are insufficient. However, these data-driven methods are highly dependent on their training data and can fail in surprising and dramatic ways with little warning. To ensure the safety of the vehicle and the surrounding environment, it is imperative that the navigation system is able to recognize the predictive uncertainty of the perception model and respond safely and effectively in the face of uncertainty. In an effort to enable safe navigation under perception uncertainty, we develop a probabilistic and reconstruction-based competency estimation (PaRCE) method to estimate the model’s level of familiarity with an input image as a whole and with specific regions in the image. We find that the overall competency score can accurately predict correctly classified, misclassified, and out-of-distribution (OOD) samples. We also confirm that the regional competency maps can accurately distinguish between familiar and unfamiliar regions across images. We then use this competency information to develop a planning and control scheme that enables effective navigation while maintaining a low probability of error. We find that the competency-aware scheme greatly reduces the number of collisions with unfamiliar obstacles, compared to a baseline controller with no competency awareness. Furthermore, the regional competency information is particularly valuable in enabling efficient navigation.
@@ I INTRODUCTION
Figure 1: We develop a probabilistic reconstruction-based method for estimating the overall and regional competency of a perception model for a given input image. This competency score and competency map are used to design a sampling-based path planning algorithm that evaluates the minimum competency associated with sampled paths, removes trajectories that fall below the minimum probability threshold, and selects the best path towards the goal. An LQR-based controller is then used to track the planned path.
@@
To address these limitations, we develop a competency-aware perception, planning, and control framework to enable safe navigation in complex environments under DNN-based perception model uncertainty (Figure 1). In particular, (1) we develop an overall competency score that represents the probability that the model’s prediction is correct for a given image. (2) We propose a regional competency method that estimates the probability that regions in the environment are unfamiliar to the model. (3) We generate a method for evaluating the probability of error associated with a proposed control sequence based on the competency estimates. (4) We develop a navigation scheme that maintains efficient performance, while ensuring the probability of error remains below a user-defined error threshold. (5) We evaluate this control scheme in a challenging lunar environment with unfamiliar obstacles and demonstrate the ability of our navigation framework to successfully and efficiently navigate to goal regions, while maintaining a low probability of error.
@@ II BACKGROUND & RELATED WORK
Field robots used in search and rescue, planetary exploration, and agricultural tasks must navigate in unstructured environments, which lack clearly viable paths or helpful landmarks [1]. In such environments, perception provides the necessary information to make the vehicle aware of the surrounding environment. Within this context, most work has focused on DNN-based methods to analyze the capability of a UGV to stably reach a terrain region and determine appropriate control actions [1]. Without properly considering the uncertainty surrounding the predictions of these DNN-based perception models, vehicles that rely on these models can face unsafe situations and system failures. This has led to a branch of uncertainty-aware navigation focused on dealing with perception model prediction uncertainty.
@@ II-A Uncertainty Quantification (UQ)
Predictive uncertainty generally encapsulates both data/aleatoric uncertainty, which arises from complexities of the data (i.e., noise, class overlap, etc.), and model/epistemic uncertainty, which reflects the ability of the perception model to capture the true underlying model of the data [8]. The modeling of these uncertainties can generally be divided into methods based on (1) Bayesian neural networks (BNNs) that extract uncertainty as a statistical measure over the output of a BNN [9, 10], (2) deterministic neural networks (NNs), often relying on Monte Carlo (MC) dropout as approximate Bayesian inference [11], and (3) ensembles of NNs that combine the predictions of multiple deterministic networks to form a probability density function [12]. While these methods address the typical overconfidence in NN predictions with better-calibrated confidence scores, they generally focus on predictions for in-distribution inputs, which come from the same distribution as the training data. These approaches are not sufficient in general to appropriately assign confidence scores for out-of-distribution (OOD) inputs that differ significantly from those seen during training [4, 13].
@@ II-D Uncertainty-Aware Navigation
There are various approaches that seek to enable short-term, local control under perception model predictive uncertainty. Existing approaches tend to rely on one of the UQ methods discussed in Section II-A, using either BNNs [48, 49], MC dropout [50], or ensembles [51, 52, 53] to estimate uncertainty and inform the control scheme. As mentioned previously, while these methods can effectively estimate some aspects of predictive uncertainty, they are insufficient for quantifying uncertainty associated with OOD samples, which a UGV is bound to encounter when navigating previously unseen or dynamic environments. There has also been work that leverages uncertainty estimation approaches that relate more closely to the distance-based OOD detection methods [54, 55] and classification-based anomaly detection methods [56, 57]. While these approaches do not fully capture the predictive uncertainty of the perception models, they are better equipped to handle inputs that differ significantly from those seen during training.
We develop a reconstruction-based method for predictive uncertainty estimation with the end goal of probabilistically safe control of a UGV. Our approach for uncertainty estimation has several benefits compared to existing methods. (1) It captures multiple aspects of uncertainty and can effectively identify OOD or anomalous samples, unlike traditional UQ techniques. (2) In contrast to methods that simply seek to identify samples that differ significantly from the training set, our method is directly tied to the predictive uncertainty of the perception model used in a control task. Furthermore, rather than making a binary decision, as is typical in OOD and anomaly detection tasks, our method is probabilistic and can be incorporated into a probabilistically safe control scheme. (3) Finally, our method can characterize uncertainty at a regional level, rather than considering uncertainty only for the image as a whole. This offers a greater amount of model explainability and provides more flexibility in designing an uncertainty-aware planning and control scheme. We show how our method for predictive uncertainty estimation can be used to evaluate the probability of error associated with vehicle trajectories and to select a preferred trajectory based on the desired level of conservativeness. [^1] The code for reproducing our methods and results is available on GitHub: https://github.com/sarapohland/parce-nav.git.
@@ IV-C Estimating Regional Model Competency
Suppose that, in addition to estimating the overall competency score for the image as a whole, we wish to estimate a regional competency score for each segmented region in the image. Now, instead of letting $\bm{X}$ be the entire input image, $\bm{X}$ is a segmented region of the image. In this work, segmented regions are determined by the Felzenszwalb segmentation algorithm [58]. We follow roughly the same procedures to estimate the probability that each region in the input image came from the same distribution as the training samples. In this approach, rather than designing an autoecoder to reconstruct the input image, we design an image inpainting model to reconstruct a missing segment of the input image and measure the average reconstruction loss over the pixels corresponding to that image segment.
@@ V COMPETENCY-AWARE NAVIGATION
We use the overall model competency estimates (§IV-A) and regional competency estimates (§IV-C) to design navigation schemes with varying levels of competency awareness. We develop a model of the vehicle dynamics (§V-A), use this model to design a competency-aware path planner (§V-B), and develop a path-tracking controller (§V-C).
@@ V-A Model of Vehicle Dynamics
Figure 4: Three examples of competency-aware planning. Column 1: Input images to perception model. Column 2: Probabilistic regional competency estimates. Blue indicates high competency (around 1.0) and red indicates low competency (around 0.0). Column 3: Sampled trajectories colored by associated minimum competency score. Column 4: Trajectories deemed safe by the competency-aware planner.
@@ V-B Competency-Aware Path Planning
We develop five planners that utilize different competency information and strategies for responding to low competency. (1) The planner that uses only overall competency selects the best trajectory in the same way as the baseline planner when overall competency is high. If the competency score falls below some probability threshold, the planner resorts to a safe action sequence–backing up slightly then turning. (2) The turning-based planner that utilizes only regional competency information checks if regions in the vicinity of the robot are associated with low perception model competency. If those regions fall below some probability threshold, the vehicle backs up slightly and turns away from low competency regions. (3) The turning-based planner that uses both overall and regional competency checks both measures of competency before resorting to a safe response. (4) The trajectory-based planner that uses only regional competency information maps sampled paths onto the regional competency map, determines the minimum competency value associated with each path, and removes those that fall below the threshold. The planner will then select the best path among those remaining. If there is no safe path, the vehicle will back up slightly and turn away from the low competency regions. (5) The trajectory-based planner that utilizes both overall and regional competency behaves similarly but only considers regional competency when overall competency is low. Three examples of this trajectory-based planner with full competency-awareness are shown in Figure 4.
@@ V-D Evaluation of Navigation Performance
We evaluate the navigation capability of the full perception, planning, and control pipeline (Figure 1) with varying levels of competency awareness and different responses to low model competency. We compare (1) a baseline controller that uses no competency information, (2) a turning-based controller that uses only overall competency scores, (3) a turning-based controller that uses only regional competency maps, (4) a trajectory-based controller that uses only regional competency information, (5) a turning-based controller that uses both overall competency scores and regional competency maps, and (6) a trajectory-based controller that uses both overall and regional competency information.
@@
TABLE I: A comparison of (1) a baseline controller with no competency awareness, (2) a turning-based controller with only overall competency information, (3) a turning-based controller with only regional competency information, (4) a trajectory-based controller with only regional competency information, (5) a turning-based controller with overall and regional competency information, and (6) a trajectory-based controller with both types of competency information.
@@ VI CONCLUSIONS
We aim to improve the safety of a UGV navigating according to a perception-based system in a complex environment with obstacles that are unfamiliar to the perception model. We believe that integrating information about the predictive uncertainty of the perception model into the navigation scheme will improve the safety and performance of the system. With this idea, we propose a novel overall competency score that can distinguish between correctly classified ID samples, misclassified ID samples, and OOD samples. We expand on this approach to develop a regional competency map that can visualize regions in an image that are familiar to the perception model and those that are not. We then explore how this competency information can be integrated into a planning and control scheme for UGV navigation.